What I the answer to 3y-5y+10=36
Found 2 solutions by Gogonati, Alan3354:
Answer by Gogonati(855)

The answer must be, solve the equation:
3y-5y+10=36, Add -10 to both sides.
3y-5y+10-10=36-10, Combine like terms.
-2y=26, Divide both sides by -2
y=-13.
Answer: The number -13 is the solution.
Check: 3(-13)-5(-13)+10=36
-39+65+10=36
-39+75=36
36=36.
Done.

Answer by Alan3354(69443)

What I the answer to 3y-5y+10=36
-------------------
What is the question?
If you want to solve for y, say that.
The other tutor did that.